# Joint Estimation of Identity Verification and Relative Pose for Partial Fingerprints

### v1

@@ V-E Alignment Accuracy
Figure 12: Estimation errors of alignment on Hybrid DB_B in image size $160\times 160$ with respective to relative rotation angle $\left|\theta\right|$ specifically, (a) rotation errors and (b) translation errors - along with relative translation difference $\left|T\right|$, which includes (c) rotation errors and (d) translation errors. Lines and corresponding light-colored bands reflect the mean and standard deviation.
@@ V-F Efficiency
In this subsection, we present a comparison of the deployment efficiency across different algorithms. As can be observed from Fig 13, our proposed method achieves a new state-of-the-art trade-off between matching performance, speed, and size. This clearly highlights our streamlined joint estimation framework of pose and identity, as well as the appropriate and efficient network design. The calculated time covers a complete process from inputting an original image pair to outputting the final estimation result. All algorithms are implemented in Python and the batch size is set to 1, which are deployed on a single NVIDIA GeForce RTX 3090 GPU with a 2.4 GHz CPU.
@@ V-G Ablation Study
Extensive ablation experiments are conducted in Table IV to examine the effects of different modules and strategies introduced in our network. Overall, the CNN based approach outperforms Vision transformer (ViT) based approach, while the CNN-Transformer hybrid series achieves the best overall performance. We attribute it to the respective advantages of convolutional and transformer blocks in extracting local features and establishing global connections. In addition, the introduction of joint estimation strategy significantly reduces the EER by 16.0%, 38.3% and 58.8% in above solutions, which proves that there is indeed an exploitable promotion relationship between pose prediction and identity verification. At the same time, the additional alignment head can be removed during deployment so that the matching speed will not be affected. Finally, the comparison between last two rows demonstrated that our proposed lightweight pre-training task can further improve the performance at a small extra training time cost.
TABLE IV: Ablation Study of the Proposed Network with Different Modules and Strategies on FVC2002 DB1_A with image size $160\times 160$.
@@
    indicates that module has been pretrained.
    each represents specific stages in proposed network, specifically ‘Encoder’ for stage 1 & 2, ‘Decoder’ for stage 3 & 4 - C, ‘Align Head’ for stage 4 - R.
@@ VI Conclusion
In this paper, we propose a joint estimation framework of identity verification and relative pose for partial fingerprints. A novel CNN-Transformer hybrid network is presented to combine their advantages in feature extraction and interaction, promoting the attention and utilization of valuable information. In addition, a lightweight pre-train task is designed to improve the representation ability of feature encoder by reconstructing enhanced images. Comprehensive experiments on multiple databases demonstrate the effectiveness and superiority of our proposed method. Futher studies will focus on further improving the performance on extremely low quality fingerprints, and extend the method to single-to-multiple scenarios.

### v2

@@ V-D Alignment Accuracy
Figure 9: Estimation errors of alignment on Hybrid DB_B in image size $160\times 160$. (a) Rotation errors and (b) Translation errors under varying $\left|\theta\right|$. (c) Rotation errors and (d) Translation errors under varying $\left|T\right|$. Lines and corresponding light-colored bands reflect the mean and standard deviation.
@@
V-E Visual Analysis
Figure 10: Representative visualization results of genuine matches based on occlusion sensitivity [21]. Red areas indicate a large influence on matching judgment, while blue areas indicate the opposite. Numbers on the bottom right are the accuracy of corresponding methods when they are verified.
In order to provide more specific and intuitive comparison with interpretability, we employ occlusion sensitivity [21] to approximate the contribution of distinct local areas to the overall matching judgment. Fig. 10 shows four representative visualization results. Query fingerprint in images pairs is centered and search fingerprint is rigidly aligned based on the ground truth as background. Subsequently, a $32\times 32$ mask is slid and applied to search fingerprints and the differences between scores before and after occlusion are calculated, which are normalized and overlaid as the heatmap of search fingerprint. Considering that there is no reference relationship between the matching scores directly obtained (or mapped) by different algorithms, we convert these scores into a comparable probability form and attached them in corresponding visualization results. Specifically, given the genuine matching score $s$ and overall distributions (shown in Fig. 7), the verification accuracy can be expressed as:
$P\left(y=1|s\right)$ (15)
where $P$ and $f$ represent the probability and probability density, $y$ represents the identity to be verified (1 is the same, 0 is different). It can be seen that algorithms based on key points [6, 9] and region representation [12, 13] expose their respective shortcomings when there is a lack of sufficient features in overlapping areas (row 1) and significant differences in ridge texture of non overlapping areas (row 2). PFVNet [7] has clear advantages in both cases, even with minimal overlap (row 3). However, it fails in the last example with confusion in rotation. At the same time, our method demonstrates optimal stability and robustness in focusing on overlapping regions and gives correct decisions with the highest accuracy.
Meanwhile, some failure cases in Fig. 11 show that current JIPNet still needs to be improved in some extreme scenarios, such as: (i) sever image defacement caused by incorrect collection conditions (row 1); (ii) significant misalignment of ridges caused by finger distortion (row 2); (iii) extremely weak texture, which refers to almost no minutiae or changes in orientation, as well as large modal differences (row 3).
@@ V-F Efficiency
In this subsection, we present a comparison of deployment efficiency across different algorithms. As shown in Fig 12, our proposed method achieves a new state-of-the-art trade-off between matching performance, speed, and size. This clearly highlights our streamlined joint estimation framework of pose and identity, as well as the appropriate and efficient network design. The calculated time covers a complete process from inputting an original image pair to outputting the final estimation result. All algorithms except for VeriFinger [9] (encapsulated SDK in C / C++) are implemented in Python and the batch size is set to 1, which are deployed on a single NVIDIA GeForce RTX 3090 GPU with an Intel Xeon E5-2680 v4 CPU @ 2.4 GHz.
@@ V-G Ablation Study
TABLE V: Ablation Study of the Proposed Network with Different Modules and Strategies on FVC2002 DB1_A with image size $160\times 160$.
@@
    indicates the corresponding module has been pretrained.
    each represents specific stages in proposed network, specifically Encoder for Stage 1 & 2, Decoder for Stage 3 & 4 - C, Align Head for Stage 4 - R.
Extensive ablation experiments are conducted to examine the effects of different modules and strategies introduced in our network. As shown in Table V, overall, the CNN based approach outperforms Vision transformer (ViT) based approach, while the CNN-Transformer hybrid series achieves the best performance. We attribute it to the respective advantages of convolutional and transformer blocks in extracting local features and establishing global connections. In addition, the introduction of parallel relative alignment estimation head achieves great success in all three types of network architectures, which proves that there is indeed an exploitable promotion relationship between pose prediction and identity verification. At the same time, the additional alignment head can be pruned during deployment so that the matching speed will not be affected. Finally, the comparison between last two rows demonstrated that our proposed lightweight pre-training task can further improve the performance at a small extra training time cost.
@@ VI Conclusion
In this paper, we propose a joint framework of identity verification and pose alignment for partial fingerprints. A novel CNN-Transformer hybrid network, named JIPNet, is presented to combine their advantages in feature extraction and information interaction, promoting the attention and utilization of valuable information. In addition, a lightweight pre-training task is designed to improve the representation ability of feature encoder by reconstructing enhanced images. Comprehensive experiments on multiple databases demonstrate the effectiveness and superiority of our proposed method. Future studies will focus on further improving the performance in more challenging scenarios such as extreme low quality and small overlap, and extending our proposed method to one-to-many matching tasks.